Find an equation of the line containing the given pair of points. (-1,-6) and (-7,-5)
Answer by rfer(16322)

m=(-5--6)/(-7--1)
m=1/-6
m=-1/6
---------------
y=mx+b
-6=-1/6*-1+b
-6=1/6+b
-6/(1/6)=b
b=-36
-----------------
y=-1/6x-36